# Gradual smoothing: strong hypercontractivity and logarithmic Sobolev inequalities

### v1

@@ Abstract.
We study the possibility of a gradual improvement as time progresses of the regularity of solutions to evolution problems of parabolic type driven by Lévy operators, not necessarily translation invariant. In the course of our analysis we study the equivalence between general smoothing effects and a family of logarithmic Sobolev inequalities. This equivalence allows us to identify a new type of regularization, strong hypercontractivity, characterized by the existence of a time at which solutions belong to every $L^{p}$ space with $p$ finite. It can also be used to prove logarithmic Sobolev inequalities in a context not previously seen in the literature. We then show that any purely nonlocal Lévy operator whose kernel is comparable to that of $\log(I-\Delta)$ is strongly hypercontractive, but fails to be supercontractive and, consequently, also fails to be ultracontractive. Finally, in the translation-invariant case, we also prove that solutions get bounded eventually.
@@ 1.1. Motivation and goal
We investigate the possibility of a gradual improvement as time evolves of the regularity of solutions to
@@
for linear operators $\mathcal{L}$ as general as possible. As part of our analysis, we prove that a general (quantitative) notion of hypercontractivity is equivalent to a family of logarithmic Sobolev inequalities. This characterization leads to the introduction of a new regularization phenomenon, which we call strong hypercontractivity.
@@
In accordance with the standard terminology in the literature, we say that the semigroups $e^{-t(-\Delta)}$ and $e^{-t(-\Delta)^{\sigma/2}}$ are ultracontractive. With a slight abuse of terminology, we will also refer to their infinitesimal generators, $\Delta$ and $(-\Delta)^{\sigma/2}$, as ultracontractive.
Let us now consider operators $\mathcal{L}_{0}$ of the form (1.1) that are: of order zero, i.e., given by a kernel $J$ integrable in the second variable; and translation invariant, i.e., with $J(x,y)=\widetilde{J}(x-y)$ for some nonnegative symmetric function $\widetilde{J}$. Let us fix $\|\widetilde{J}\|_{1}=1$. These operators do not define an ultracontractive semigroup. Indeed, when  $\mathcal{L}=\mathcal{L}_{0}$, the solution to ($\textup{P}_{\mathcal{L}}$) is given by $u(t)=e^{-t}u_{0}+v(t)*u_{0}$ for some smooth function $v$, the so-called regular part of the heat kernel; see for instance [14]. Therefore, $u(t)$ is only as good as the initial datum is; it does not gain integrability nor regularity.
@@
(1.3) $\widehat{H_{t}}(\xi)=e^{-tm(\xi)}=(1+|\xi|^{2})^{-t}\in L^{1}(\mathbb{R}^{N})\quad\textup{if and only if }t>N/2.$
@@
(1.4) $\|e^{-t\mathcal{L}}u_{0}\|_{\infty}\leq A_{p}(t)\|u_{0}\|_{p}\quad\text{for }p\geq 1,\,t>t_{\infty}=\frac{N}{2p},$
for some function $A_{p}(t)\geq 0$. Up to time $t_{\infty}$ we have hypercontractivity: a gradual gain of integrability. Furthermore, we will prove that $\widehat{H}_{t}(\xi)\in L^{\frac{N}{N-2t},\infty}(\mathbb{R}^{N})$ if $t<N/2$, whence, by weak Young’s inequality,
(1.5) $\|e^{-t\mathcal{L}}u_{0}\|_{\frac{Np}{N-2pt}}\leq A_{p}(t)\|u_{0}\|_{p}\quad\text{ for }p>1,\,0<t<\frac{N}{2p},$
so that any solution with initial data in $L^{p}(\mathbb{R}^{N})$$p>1$, is at time $t=t_{\infty}$ in every $L^{q}(\mathbb{R}^{N})$$q\in[p,\infty)$. We will refer to this property as strong hypercontractivity. Despite its close connection to eventual ultracontractivity, this notion does not necessarily entail the latter for general operators. When the gain of integrability arises instantaneously, it is referred to as supercontractivity. We remark that solutions $u(t)=H_{t}*u_{0}$ eventually become arbitrarily regular, not merely increasingly integrable.
Let us mention that solutions become as regular as desired, not just more integrable, for $t$ big enough; see Section 4.3.
@@
Singular kernels. Observe that the infinitesimal generators $(-\Delta)^{\sigma/2}$,  $\sigma\in(0,2)$, require a certain amount of regularity of a function $f$ in order for $(-\Delta)^{\sigma/2}f$ to be well defined. Indeed, the operator $(-\Delta)^{\sigma/2}$ can be written in the form (1.1), with kernel
@@
The operator $\log(I-\Delta)$ falls into the class of nonlocal operators of type (1.1) with a kernel that behaves like $|x-y|^{-N}$ when $x$ and $y$ are close. This kernel behaviour places $\log(I-\Delta)$ in an intermediate position between the two previous cases, precisely at the threshold of integrability: $|x|^{-N-\delta}$ is locally integrable if $\delta<0$, but fails to be so when $\delta>0$. We will see that all operators of the form (1.1) whose kernels $J(x,y)$ behave as $|x-y|^{-N}$ share similar smoothing properties with $\log(I-\Delta)$.
@@
Sobolev and logarithmic Sobolev inequalities. We define the Dirichlet form $\mathcal{E}(\cdot,\cdot)$ associated to operators of the form (1.1) by
@@
However, one of the most important points of the present paper is the following: while this inequality alone is enough to characterize hypercontractivity (although not in its full extent), it is not enough to understand strong hypercontractivity. To have such property, a family of pp–logarithmic Sobolev inequalities of the form
(1.11) $\||f|^{p}\log(|f|^{p})\|_{1}-\|f\|_{p}^{p}\log(\|f\|_{p}^{p})\leq C(p)\|f\|_{p}^{p}+D(p)\mathcal{E}(|f|^{p-2}f,f),$
is required to hold for all $p\geq 2$. Our first aim is to study the equivalence between general hypercontractivity and the existence of a family of logarithmic Sobolev inequalities of the above type, and deduce from this characterization logarithmic Sobolev inequalities for the operator $\log(I-\Delta)$.
@@ 1.2. Main results
We start our work by defining the concepts of entropy, hypercontractivity and pp–logarithmic Sobolev inequalities. We next prove Theorem 3.11, showing that hypercontractivity for some $p\geq 2$ implies a corresponding pp–logarithmic Sobolev inequality. In Theorem 3.12 we establish the reciprocal statement, thus proving the equivalence between hypercontractivity and logarithmic Sobolev inequalities. These two statements are, in essence, Gross’s results, suitably reformulated for our framework. Depending on the precise behaviour as $p\to\infty$ of the coefficients in the family of pp–logarithmic Sobolev inequalities, we characterize strong hypercontractivity and provide sufficient conditions for eventual ultracontractivity. Through a duality argument, we extend hypercontractivity to the range $p\in(1,2)$ in Section 3. Our approach also leads to supercontractivity and ultracontractivity results. Indeed, as a consequence of our equivalence theorems, we obtain a full characterization of supercontractivity and recover the classical sufficient conditions for ultracontractivity established in the literature. Fully understanding these equivalences represents the first main contribution of this work.
@@
To exploit the previous results, we undertake a detailed analysis of the near-zero order operator $\log(I-\Delta)$, which serves as an ideal model for the theorems above. In Section 4.3 we show that solutions to problem ($\textup{P}_{\mathcal{L}}$) involving this logarithmic operator gain regularity gradually: as time evolves, $u(t)$ enters into increasingly more regular Sobolev-type spaces. By Sobolev embeddings, this improvement yields arbitrarily high regularity for sufficiently large times and, in particular, eventual ultracontractivity. A finer, quantitative study of hypercontractivity is carried out in Theorem 4.7, which essentially reduces to a careful examination of the Sobolev embeddings associated with the Bessel potential. Building on this groundwork, we then apply the equivalence results established earlier to prove in Section 4.5 the logarithmic Sobolev inequalities satisfied by $\log(I-\Delta)$.
@@
Although the operator $\log(I-\Delta)$ has been previously considered in the literature, we have not found the associated logarithmic Sobolev inequalities obtained here stated anywhere else. To the best of our knowledge, they constitute the first example of essential logarithmic Sobolev inequalities in an infinite measure space, in the sense that our operator does not satisfy any stronger Sobolev inequality of the form (1.8). These inequalities represent the second main contribution of the paper.
@@
Then, we examine general operators whose kernels satisfy $J(x,y)\sim|x-y|^{-N}$ for $x\sim y$. We prove that such operators enjoy strong hypercontractivity, but fail to be supercontractive —and hence also ultracontractive. The argument relies on several of the results established earlier and shows that the obtained conclusion is optimal. This smoothing effect for Lévy operators that may be non-translation-invariant constitutes the third main contribution of the paper.
Finally, we turn our attention to translation-invariant operators, whose kernels take the form $J(x,y)=\widetilde{J}(x-y)$, with $\widetilde{J}$ a symmetric function. We show in Theorem 6.6 that if $\widetilde{J}\geq c\widetilde{J}_{\log}$ for some $c>0$, at least for small interactions, then solutions to ($\textup{P}_{\mathcal{L}}$) for $\mathcal{L}$ defined by the kernel $\widetilde{J}$ enter into increasingly more regular Sobolev-type spaces just like the corresponding solutions for $\log(I-\Delta)$. We then show that, in general, the lower bound $\widetilde{J}(z)\geq c|z|^{-N}$ for some $c>0$ and $z\sim 0$ is precisely the condition ensuring the regularity improvement. In particular, this requirement characterizes eventual ultracontractivity and the gain in regularity of the fundamental solution; see Theorem 6.7.
@@ 1.3. Precedents
where the left-hand side of the inequality is the so-called entropy of $f$. This is the first instance in which the equivalence between hypercontractivity and a logarithmic Sobolev inequality is established.
That work already includes what we refer to here as pp–logarithmic Sobolev inequalities. When an operator $\mathcal{L}$ satisfies a pp–logarithmic Sobolev inequality, Gross refers to it as a Sobolev generator of index $p$. However, in his paper, these inequalities are used only as an intermediate technical tool for establishing Nelson’s hypercontractivity, and the fundamental importance of this entire family of inequalities has until now gone unrecognized; see Section 3 for further details.
@@
The contractivity properties of subordinate semigroups, particularly in contexts involving infinite measures such as Lebesgue’s measure, have been extensively analyzed in the literature. Subordinate semigroups arise as families of contractions generated by applying suitable functions to a fixed operator, and one seeks to identify structural conditions on these functions that guarantee different forms of contractivity. A systematic treatment of eventual ultracontractivity within this framework was first established in [6, Theorem 3.1]. In this theory, the operator $\log(I-\Delta)$ appears naturally as the functional calculus realization of the mapping $s\mapsto\log(1+s)$ applied to the positive operator $\Delta$.
A related line of research aims at establishing connections between conditions on the subordination functions and functional inequalities of Nash or super-Poincaré type, like the ones obtained in [26]. These inequalities arise naturally within the framework of Bernstein functional calculus, see for instance the monograph [42]. The stability of such inequalities under subordination is proved in [26]. They are then used to characterize hypercontractivity in [43, Proposition 13]. Nevertheless, whilst there is a direct relation between Nash, super-Poincaré inequalities and ultracontractivity, it seems that they are not sufficient if one wants to prove eventual ultracontractivity; see Section 7 for a more detailed explanation.
@@
One of the main sources of inspiration for the present work is [11], where it is shown that, for translation-invariant Lévy operators, the threshold for ultracontractivity occurs precisely when the function $\widetilde{J}$ defining the kernel satisfies $\widetilde{J}(z)\sim|z|^{-N}$ near the origin. Further related developments include the analysis in [19] of the energy space associated with near-zero Lévy operators yielding, among other results, a Hardy inequality and an embedding into a Lorentz space; and the proofs in [31, 33] that solutions to a related elliptic problem with Dirichlet conditions are bounded and continuous.

### v2

@@ Abstract.
We study the possibility of a gradual improvement as time progresses of the regularity of solutions to evolution problems of parabolic type driven by Lévy-type operators, not necessarily translation invariant. In the course of our analysis we study the equivalence between general smoothing effects and a family of logarithmic Sobolev inequalities. This equivalence allows us to identify a new type of regularization, strong hypercontractivity, characterized by the existence of a time at which solutions belong to every $L^{p}$ space with $p$ finite. It can also be used to prove logarithmic Sobolev inequalities in a context not previously seen in the literature. We then show that any purely nonlocal Lévy-type operator whose kernel is comparable to that of $\log(I-\Delta)$ is strongly hypercontractive, but fails to be supercontractive and, consequently, also fails to be ultracontractive. Furthermore, in the translation-invariant case, we also prove that solutions get bounded eventually and start improving in differentiability right after doing so. Finally, we show that this behaviour only appears if the kernel defining the operator behaves as $|x-y|^{-N}$ for small interactions ($0^{+}$-order operators): more singular kernels yield instantaneous smoothing, while less singular ones do not produce any regularization.
@@ 1.1. Motivation and goal
We investigate the possibility of a gradual improvement, as time evolves, of the regularity of solutions to
@@
for linear operators $\mathcal{L}$ as general as possible. By gradual smoothing of solutions we mean that solutions belong to progressively better integrability spaces as time evolves (hypercontractivity), and once they become bounded, their differentiability also starts improving gradually. We will study the whole improving process, although we focus more on the first part of it: the gradual improvement in integrability. As part of our analysis, we prove that a general (quantitative) notion of hypercontractivity is equivalent to a family of logarithmic Sobolev inequalities. This characterization leads to the introduction of a new regularization phenomenon, which we call strong hypercontractivity.
@@
Following the standard usage in the literature, we say that the semigroups $e^{-t(-\Delta)}$ and $e^{-t(-\Delta)^{\sigma/2}}$ are ultracontractive. With a slight abuse of terminology, we will also refer to their infinitesimal generators, $\Delta$ and $(-\Delta)^{\sigma/2}$, as ultracontractive.
Let us now consider operators $\mathcal{L}_{0}$ of the form (1.1) that are: of order zero, i.e., given by a kernel $J$ integrable in the second variable; and translation invariant, i.e., with $J(x,y)=\widetilde{J}(x-y)$ for some nonnegative symmetric function $\widetilde{J}$. Let us fix $\|\widetilde{J}\|_{1}=1$. These operators do not define an ultracontractive semigroup. Indeed, when $\mathcal{L}=\mathcal{L}_{0}$, the solution to ($\textup{P}_{\mathcal{L}}$) is given by $u(t)=e^{-t}u_{0}+v(t)*u_{0}$ for some function $v$, the so-called regular part of the heat kernel; see for instance [14]. Therefore, $u(t)$ is only as good as the initial datum is; it does not gain integrability nor regularity.
@@
(1.3) $\widehat{H_{t}}(\xi)=e^{-t\log(1+|\xi|^{2})}=(1+|\xi|^{2})^{-t}\in L^{1}(\mathbb{R}^{N})\quad\textup{if and only if }t>N/2.$
@@
(1.4) $\|e^{-t\log(I-\Delta)}u_{0}\|_{\infty}\leq A_{p}(t)\|u_{0}\|_{p}\quad\text{for }p\geq 1,\,t>t_{\infty}=\frac{N}{2p},$
for some function $A_{p}(t)>0$. Up to time $t_{\infty}$ we have hypercontractivity: a gradual gain of integrability. Furthermore, we will prove that $H_{t}\in L^{\frac{N}{N-2t},\infty}(\mathbb{R}^{N})$ if $t<N/2$, whence, by weak Young’s inequality,
(1.5) $\|e^{-t\log(I-\Delta)}u_{0}\|_{\frac{Np}{N-2pt}}\leq A_{p}(t)\|u_{0}\|_{p}\quad\text{ for }p>1,\,0<t<\frac{N}{2p},$
so that any solution with initial data in $L^{p}(\mathbb{R}^{N})$$p>1$, is at time $t=t_{\infty}$ in every $L^{q}(\mathbb{R}^{N})$$q\in[p,\infty)$. We will refer to this property as strong hypercontractivity. Despite its close connection to eventual ultracontractivity, this notion does not necessarily entail the latter for general operators. When the gain of integrability arises instantaneously, it is referred to as supercontractivity. Let us mention that solutions become as regular as desired, not just more integrable, for $t$ big enough; see Section 4.3.
@@
Singular kernels. Observe that the infinitesimal generators $(-\Delta)^{\sigma/2}$$\sigma\in(0,2)$, require a certain amount of regularity of a function $f$ in order for $(-\Delta)^{\sigma/2}f$ to be well defined. Indeed, the operator $(-\Delta)^{\sigma/2}$ can be written in the form (1.1), with kernel
@@
The operator $\log(I-\Delta)$ falls into the class of nonlocal operators of type (1.1) with a kernel that behaves like $|x-y|^{-N}$ when $x$ and $y$ are close. This places it in an intermediate position between the two previous cases, precisely at the threshold of integrability: $|x|^{-N-\delta}$ is locally integrable if $\delta<0$, but not when $\delta>0$. We will show that all operators of the form (1.1) whose kernels have the same critical singularity share smoothing properties analogous to those of $\log(I-\Delta)$. This motivates the terminology $0^{+}$-order operators to denote them; in the literature, they are also referred to as near-zero order or geometrically stable operators.
@@
Sobolev and logarithmic Sobolev inequalities. We define the Dirichlet form $\mathcal{E}(\cdot,\cdot)$ associated to an operator of the form (1.1) by
@@
However, one of the main insights of this paper is that the above inequality, while sufficient to characterize a certain form of hypercontractivity, as shown by Gross [33], falls short of describing strong hypercontractivity. To recover such property, a family of pp–logarithmic Sobolev inequalities of the form
(1.11) $\||f|^{p}\log(|f|^{p})\|_{1}-\|f\|_{p}^{p}\log(\|f\|_{p}^{p})\leq C(p)\|f\|_{p}^{p}+D(p)\mathcal{E}(|f|^{p-2}f,f)$
is required to hold for all $p\geq 2$. The asymptotic behaviour of $C(p)$ and $D(p)$ for $p\to\infty$ is what will matter for strong hypercontractivity to appear.
There is no direct mechanism for transferring hypercontractivity from one operator to another. Logarithmic Sobolev inequalities, however, are energy estimates, and operators of the form (1.1) with comparable kernels satisfy the same such estimates. Our first aim is to show that general hypercontractivity is equivalent to the existence of an appropriate family of logarithmic Sobolev inequalities of the above type. This reduces the problem: hypercontractivity for any $0^{+}$-order operator follows once it is proved for a single model in the class. We therefore focus on $\log(I-\Delta)$.
@@ 1.2. Main results
We start our work by defining the concepts of entropy, hypercontractivity and pp–logarithmic Sobolev inequalities. We next prove Theorem 3.11, showing that hypercontractivity for some $p\geq 2$ implies a corresponding pp–logarithmic Sobolev inequality. In Theorem 3.12 we establish the reciprocal statement, thus proving the equivalence between hypercontractivity and logarithmic Sobolev inequalities. These two statements are, in essence, Gross’s results, suitably reformulated for our framework. Depending on the precise behaviour as $p\to\infty$ of the coefficients in the family of pp–logarithmic Sobolev inequalities, we characterize strong hypercontractivity and provide sufficient conditions for eventual ultracontractivity. Through a duality argument, we extend hypercontractivity to the range $p\in(1,2)$ in Section 3. Our approach also leads to supercontractivity and ultracontractivity results. Indeed, as a consequence of our equivalence theorems, we obtain a full characterization of supercontractivity and recover the classical sufficient conditions for ultracontractivity established in the literature. Fully understanding these equivalences between smoothing effects and logarithmic Sobolev inequalities represents the first main contribution of this work.
@@
In order to exploit the previous results, we undertake a detailed analysis of the $0^{+}$-order operator $\log(I-\Delta)$, which serves as an ideal model for the theorems above. In Section 4.3 we show that solutions to problem ($\textup{P}_{\mathcal{L}}$) involving this logarithmic operator gain regularity gradually: as time evolves, $u(t)$ enters into increasingly more regular Sobolev-type spaces. By Sobolev embeddings, this improvement yields arbitrarily high regularity for sufficiently large times and, in particular, eventual ultracontractivity. A finer, quantitative study of hypercontractivity is carried out in Theorem 4.7, which essentially reduces to a careful examination of the Sobolev embeddings associated with the Bessel potential $(I-\Delta)^{-t}$ as $t\to 0$. Building on this groundwork, we then apply the equivalence results established earlier to prove in Section 4.5 the logarithmic Sobolev inequalities satisfied by $\log(I-\Delta)$.
@@
Although the operator $\log(I-\Delta)$ has been previously considered in the literature, we have not found the associated logarithmic Sobolev inequalities obtained here stated anywhere else. To the best of our knowledge, they constitute the first example of essential logarithmic Sobolev inequalities in an infinite measure space, in the sense that our operator does not satisfy any stronger Sobolev inequality of the form (1.8). Furthermore, the logarithmic Sobolev inequality serves a different purpose here than in the previous literature. It is typically used to obtain asymptotic results via entropy methods; in our setting, it provides the best energy estimate (Sobolev embedding) expected for $0^{+}$-order operators. Establishing these inequalities is the second main contribution of the paper.
@@
Then, we examine general operators whose kernels satisfy $J(x,y)\sim|x-y|^{-N}$ for $x\sim y$. We prove that such operators enjoy strong hypercontractivity, but fail to be supercontractive, and hence also ultracontractive. The proof relies essentially on the equivalence established earlier, together with a fine analysis of the model operator $\log(I-\Delta)$. The key idea is to transfer hypercontractivity via logarithmic Sobolev inequalities for operators with comparable kernels. We also show that the result is optimal: regularization is necessarily gradual. This smoothing effect for Lévy operators that need not be translation invariant is the third main contribution of the paper.
Finally, we turn our attention to translation-invariant operators, whose kernels take the form $J(x,y)=\widetilde{J}(x-y)$, with $\widetilde{J}$ a symmetric function. We show in Theorem 6.6 that if $\widetilde{J}(z)\geq c|z|^{-N}$ for some $c>0$ and $|z|\ll 1$, then solutions to ($\textup{P}_{\mathcal{L}}$) for $\mathcal{L}$ defined by the kernel $\widetilde{J}$ enter into increasingly more regular Sobolev-type spaces, just like the corresponding solutions for $\log(I-\Delta)$. Thus, solutions improve first in integrability and then in differentiability after getting bounded. Finally, we show in Theorem 6.7 that this gradual smoothing phenomenon only appears if the kernel defining the operator behaves as $|x-y|^{-N}$ for small interactions. Any radial perturbation making the kernel slightly more or less singular breaks this behaviour: if $\ell(|z|)=\widetilde{J}(|z|)|z|^{N}$ verifies $\lim_{z\to 0}\ell(z)=\infty$, solutions get bounded and $\mathcal{C}^{\infty}$ instantaneously, just as when $\mathcal{L}=(-\Delta)^{\sigma/2}$ for $\sigma\in(0,2]$; and if $\lim_{z\to 0}\ell(z)=0$, solutions remain only as regular as the initial data, just as when $\mathcal{L}$ is defined via an integrable kernel. These results in the translation-invariant setting constitute the fourth and final main contribution of the paper.
@@ 1.3. Precedents
where the left-hand side is the so-called entropy of $f$. This work was the first to establish the equivalence between hypercontractivity and a logarithmic Sobolev inequality.
That paper already includes what we denote here as pp–logarithmic Sobolev inequalities. When an operator $\mathcal{L}$ satisfies a pp–logarithmic Sobolev inequality, Gross refers to it as a Sobolev generator of index $p$. However, in his work, these inequalities are used only as an intermediate technical tool for establishing Nelson’s hypercontractivity, and the fundamental importance of this entire family of inequalities has until now gone unrecognized; see Section 3 for further details.
@@
The contractivity properties of subordinate semigroups, particularly in contexts involving infinite measures such as Lebesgue’s measure, have been extensively analysed in the literature. Such semigroups arise naturally within the framework of Bernstein functional calculus, see for instance the monograph [48]. A systematic treatment of eventual ultracontractivity within this framework was first established in [6, Theorem 3.1]. In this theory, the operator $\log(I-\Delta)$ appears naturally as the functional calculus realization of the mapping $s\mapsto\log(1+s)$ applied to the positive operator $\Delta$. This theory also allows to establish functional inequalities of Nash or super-Poincaré type, like the ones obtained in [29]. They are then used to characterize hypercontractivity in [49, Proposition 13]. Nevertheless, whilst there is a direct relation between Nash, super-Poincaré inequalities and ultracontractivity, it seems that they are not sufficient if one wants to prove eventual ultracontractivity; see Section 7 for a more detailed explanation.
@@
One of the main sources of inspiration for the present work is [11], where it is shown that, for translation-invariant Lévy operators, the threshold for ultracontractivity occurs precisely when the function $\widetilde{J}$ defining the kernel satisfies $\widetilde{J}(z)\sim|z|^{-N}$ near the origin. Further related developments include the analysis in [19] of the energy space associated with $0^{+}$-order Lévy operators yielding, among other results, a Hardy inequality and an embedding into a Lorentz space; and the proofs in [35, 37] that solutions to a related elliptic problem with Dirichlet conditions are bounded and continuous.
Finally, it is worth mentioning that the operator $\log(I-\Delta)$ can be understood as the following limit of fractional relativistic Schrödinger operators $(I-\Delta)^{\alpha}-I$,
These operators are well known and widely studied; see, for example, [26, 27, 28, 34, 40, 44] for a representative (though not exhaustive) list of references. Observe that $\log(I-\Delta)$ can also be seen as the derivative at $\alpha=0$ of the Bessel operators $(I-\Delta)^{\alpha}$. We do not make use of these two limits to obtain information about $\log(I-\Delta)$. However, what will be of importance is that $\log(I-\Delta)$ is the infinitesimal generator associated to the Bessel potential $(I-\Delta)^{-t}$ semigroup.